OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4

OpenStudy (anonymous):

@amistre64 @satellite73 @thomaster @touseii45 @TheHero

OpenStudy (anonymous):

use your point-slope formula

OpenStudy (anonymous):

i dont know it can show me

OpenStudy (anonymous):

\[y - y1=m(x-x1)\]

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

well you point (1, -5) is your x1 and y1 so plug that in

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and the y=2x+4 is in slope intercept form which is y=mx+b meaningthat your 2 is the m in the point-slope formula

OpenStudy (anonymous):

Are you with me so far?

OpenStudy (anonymous):

@Helpmeplease99